OpenStudy (anonymous):

Factor completely: 8x^2+20x-12

OpenStudy (anonymous):

=4(2x^2 + 5x - 3) = 4 (2x - 1 )( x + 3)

OpenStudy (anonymous):

these need practice before you become confident - do u need explanation?

OpenStudy (anonymous):

I'm still a little confused about the steps. This is my first time working with these kinds of problems.

OpenStudy (anonymous):

ok first take out any common factors - in this case all the numbers have a factor 4

OpenStudy (anonymous):

And then you divide each number by 4?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Don't worry about the 4, after you factor it out just disregard it

OpenStudy (anonymous):

Worry about what is left after taking out the greatest common factor

OpenStudy (anonymous):

When you do that, you are left with the 2^2 + 5x-3

OpenStudy (anonymous):

From there, make two binomials

OpenStudy (anonymous):

that's why the 2^2 is no longer squared, because you split the x's

OpenStudy (anonymous):

now - if you make 3 columns as follows it makes it a bit easier 2x^2 -3 +5x ----------------------------------------- 2 x 1 - 3 x 1 2 x -3 + 1 x 1 = -5 - not +5 +3 x -1 2 x +3 + -1 x 1 = +5 correct so the the 2x multiplies +3 and x multiplies -1 so we have (2x -1)(x + 3) note its first in bracket 1 x by last in bracket 2 (two outsides} and last x first 2 (two insides}

OpenStudy (anonymous):

do you get what he said?

OpenStudy (anonymous):

So, when factoring 2x^3+3x^2-32x-48